OpenStudy (anonymous):

i need HELP solving this equation!!! 5^a 1^a/4 - 2/2^a = 0

myininaya (myininaya):

\[\frac{5^a1^a}{4}-\frac{2}{2^a}=0\] is this right?

OpenStudy (anonymous):

yes

myininaya (myininaya):

1^a=1 so we have \[\frac{5^a}{4}-\frac{2}{2^a}=0\]

OpenStudy (anonymous):

ok

myininaya (myininaya):

\[\frac{5^a}{4}=\frac{2}{2^a}\] cross multiply \[5^a*2^a=2*4\] \[(5*2)^a=8\] \[10^a=8\] take log_10 on both sides (and i will just call it log) \[\log10^a=\log8\] \[alog10=\log8\] \[a=\frac{\log8}{\log10}\]